# UniBias: Unveiling and Mitigating LLM Bias through Internal Attention and FFN Manipulation

### v1

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in various tasks using the in-context learning (ICL) paradigm. However, their effectiveness is often compromised by inherent bias, leading to prompt brittleness—sensitivity to design settings such as example selection, order, and prompt formatting. Previous studies have addressed LLM bias through external adjustment of model outputs, but the internal mechanisms that lead to such bias remain unexplored. Our work delves into these mechanisms, particularly investigating how feedforward neural networks (FFNs) and attention heads result in the bias of LLMs. By Interpreting the contribution of individual FFN vectors and attention heads, we identify the biased LLM components that skew LLMs’ prediction toward specific labels. To mitigate these biases, we introduce UniBias, an inference-only method that effectively identifies and eliminates biased FFN vectors and attention heads. Extensive experiments across 12 NLP datasets demonstrate that UniBias significantly enhances ICL performance and alleviates prompt brittleness of LLMs.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional capabilities in various natural language processing (NLP) tasks, employing the in-context learning (ICL) paradigm. This paradigm conditions LLMs on a context prompt comprising of a few example-label pairs [1, 5].
@@
Despite their impressive performance, LLMs are prone to prompt brittleness, characterized by high sensitivity to the choice [29] and order [14] of examples, and prompt formatting [15], as demonstrated in Figure 1. Such prompt brittleness is found to be arise from the bias in LLMs towards predicting certain answers [29]. The presence of the bias undermines the robustness and adaptability of LLMs in diverse applications.
@@
Extensive research has focused on identifying factors that lead to LLM bias and strategies for mitigation. For instance, vanilla label bias [7] and recency bias [29] demonstrate the LLM’s inherent non-contextual preference for certain labels and contextual preference for specific positions, respectively. Additionally, several calibration methods [7, 10, 29] are proposed to counteract the bias by adjusting decision boundaries of model output probabilities. However, these approaches are derived from external observations or adjustments of LLM outputs, leaving the internal mechanisms within LLMs that cause such bias poorly understood.
@@
In this work, we investigate the internal mechanism of LLM bias, specifically how feedforward neural networks (FFNs) and attention heads contribute to such bias. Building on findings in mechanistic interpretability [6, 4], we assess the contribution of individual attention heads and FFN vectors [^1] FFN vector refers to the value vector in the second weight matrix of the FFN layer. We elaborate on this in Section 2.1 to label predictions in LLMs. By identifying FFN vectors and attention heads that convey biased influences towards label prediction, we reveal the internal mechanisms behind several key bias factors, including vanilla label bias [7], recency bias [29], and selection bias [30]. For instance, our analysis of FFN vectors without input context demonstrates that their cumulative impact biases the LLM towards specific labels, indicating a non-contextual preference for certain labels, i.e., vanilla label bias. We elaborate on the background of mechanistic interpretability in Section 2.1 and present our findings on the internal mechanisms of LLM biases in next section.
@@
we pose the question: Can we identify the biased components of LLMs and mitigate their detrimental impact on label prediction? Motivated by this intuition, we propose UniBias, an inference-only method designed to identify and eliminate biased FFN vectors and attention heads in LLMs. Specifically, we begin by projecting each FFN vector and attention head into the vocabulary space to interpret the information conveyed by their outputs. We then detect biased components based on three criteria we defined: the relatedness criterion, the bias criterion, and the low variance criterion. After identification, we mitigate their impact by masking these biased components. Extensive experimental results demonstrate that LLMs, from which biased components have been removed, consistently outperform their original counterparts by a significant margin. Further, as illustrated in Figure 1, our method significantly improves both the performance and robustness of ICL with perturbations of various design settings.
@@
Figure 1: illustrates the prompt brittleness of ICL and the effectiveness of our method in mitigating this issue. Experiments are conducted in one-shot setting, using SST2 [20] dataset for experiments on example selection and prompt formatting and AGnews [28] dataset for example order experiment due to more diverse combination of orders.
@@
  • In contrast to existing works based on external observation of LLM outputs, we unveil the internal mechanisms within LLMs that lead to their bias towards predicting certain answers.
  • We propose the UniBias method to mitigate LLM bias by identifying and eliminating biased FFN vectors and attention heads within LLMs. Moreover, our method demonstrate an effective way to manipulate internal structures of LLMs.
  • Extensive experiments across 12 NLP datasets demonstrate that, by removing the biased components, our UniBias method significantly enhances ICL performance compared to the original LLM, achieving state-of-the-art results.
@@ 2.1 Background
The theoretical background of this work is based on research on mechanistic interpretability [6, 23, 8], which aims to explain the internal processes in language models (LMs), facilitating the interpretation of the contributions of individual model components to the final prediction.
@@
The Residual Stream We interpret Transformers following the view of residual stream [6, 4]. Due to the residdual connection of Transformers, each layer takes a hidden state as input, and adds information obtained by its attention layer and FFN layer to the hidden state through residual connection. In this sence, the hidden state is a residual stream passed along layers, and each attention layer and FFN layer contribute to the final prediction by adding information to the residual stream.
Attention Heads   Following Elhage et al. [6], Dar et al. [4], the output of each attention layer of LM can be computed as the sum of all its attention heads. Specifically, for $l$-th layer, the input is ${X^{l}}\in\mathbb{R}^{N\times d}$, and the attention layer is parameterized by four matrices $W_{Q}^{l}$, $W_{K}^{l}$, $W_{V}^{l}$, $W_{O}^{l}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and the rows of the output matrix can be split into $H$ parts: $W_{Q}^{\ell,j},W_{K}^{\ell,j},W_{V}^{\ell,j}\in\mathbb{R}^{d\times\frac{d}{H}}$ and $W_{O}^{\ell,j}\in\mathbb{R}^{\frac{d}{H}\times d}$, where $H$ is the number of attention heads. We then find that:
@@
FFN   In line with Geva et al. [8, 9], transformer FFN layers can be cast as linear combination of vectors. Specifically, for an input vector $\mathbf{x}^{\ell}\in\mathbb{R}^{d}$, FFN parameter matrices $\mathbf{K}^{\ell},\mathbf{V}^{\ell}\in\mathbb{R}^{d_{m}\times d}$, the FFN output can be derived as:
@@
Logit Lens   The logit lens [16] is a technique that directly decode hidden states into the vocabulary space using the unembedding matrix of the LLM for interpretation. This approach has been validated in various studies as an efficient method for interpreting the weight matrix or hidden states of LLMs [4, 11, 27, 8].
@@ 2.2 Internal Mechanisms of Bias Factors
We delve into the mechanisms behind several bias factors, analyzing the contributions of attention heads and FFN vectors to the biased predictions in LLMs. We explore vanilla label bias, position bias, and selection bias using the Llama-2 7B model [21].
Vanilla Label Bias   The vanilla label bias [7], also known as common token bias [29], is the inherent, uncontextual preference of the model towards predicting certain label names. Given the contextual nature of attention layers, our investigation focuses on the FFN layers, where we identified a corresponding uncontextual preference. Specifically, by projecting the FFN value vectors into the vocabulary space, we compute the logits for various label names for each FFN vector. Utilizing the residual stream insight, we then aggregate these logits for all FFN vectors whose label logits rank within the top $10$ over the vocabulary, reflecting uncontextual influences of FFN vectors that are effective in label prediction. This process yields what we term uncontextual accumulated FFN logits, revealing the intrinsic bias of the LLM towards predicting label names without the influence of input.
@@
Recency Bias   Recency bias refers to the tendency of LLMs to favor the label of the example at the end of the prompt [29]. By examining the behavior of attention heads within LLMs, we observe that specific heads consistently prioritize the example at the end of the prompt, providing an internal perspective on the origin of recency bias.
@@
Selection Bias   The selection bias refers that LLMs prefer to select specific option ID (like "Option A") as answers for multiple choice questions [30]. We have identified both FFN vectors and attention heads that consistently favor a specific option regardless of the ground truth label of the test sample, revealing the internal mechanism of selection bias.
We evaluate the Llama-2 7B model on the ARC dataset, which contains four options (A, B, C, D). We use a zero-shot setting to avoid the influence of position bias from multiple examples. More details are provided in Appendix A.
Experimental results are illustrated in Figure 4. Firstly, we observe that the LLM exhibits a vanilla label bias favoring option "A", as shown in the first subfigure. Additionally, we identify a biased attention head that demonstrates a position bias consistently favoring the first option regardless of the ground truth labels of the test samples (second subfigure) or changes in the sequence of options (third subfigure). Since option A is usually the first option, these two biases both lead to the LLM’s preference for option A.
@@ 3.1 Biased FFN Vectors Identification
We use $p$ unlabeled samples from the task to assess the three criteria we defined. The coefficients and label logits of an FFN vector for these samples are denoted as $\mathbf{m}=[m_{0},m_{1},\ldots,\ m_{p-1}]$ and $\mathbf{G}=[\mathbf{g}^{(0)},\mathbf{g}^{(1)},\ldots,\mathbf{g}^{(p-1)}]^{\top}\in\mathbb{R}^{p\times c}$, respectively. An FFN vector is considered biased if it meets the following conditions:
$&\frac{1}{p}\sum_{k=0}^{p-1}\text{Sum}\left(\mathbf{G}_{k,:}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\text{Sum}\left(\mathbf{g^{(k)}}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\sum_{j=0}^{c-1}g_{j}^{(k)}>th_{FFN}^{1}\\ &\frac{1}{p}\sum_{k=0}^{p-1}\text{Bias}\left(\mathbf{G}_{k,:}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\text{Bias}\left(\mathbf{g^{(k)}}\right)=\frac{1}{p}\frac{1}{c}\sum_{k=0}^{p-1}\sum_{j=0}^{c-1}\left(g_{j}^{(k)}-\mu(\mathbf{g^{(k)}})\right)>th_{FFN}^{2}\\ &CV\left(\mathbf{m}\right)=\frac{\sigma(\mathbf{m})}{\mu(\mathbf{m})}=\frac{\sqrt{\frac{1}{p}\sum_{j=0}^{p-1}\left(m_{k}-\mu(\mathbf{m})\right)^{2}}}{\frac{1}{p}\sum_{k=0}^{p-1}m_{k}}<th_{FFN}^{3}\end{aligned}\right.$
@@
The relatedness criterion is measured by the sum of label logits. The biased criterion is measured by the logit difference between each label logit and the average label logit. The low variance criterion is measured by the coefficient variance (CV) of the FFN vector coefficient across different samples, which is the standard deviation normalized by the mean of the data.
@@ 4.1 Experimental Setup
Baselines   In addition to the standard ICL, we compare our proposed UniBias with state-of-the-art LLM debiasing and calibration baselines, including Contextual Calibration (CC) [29], Domain-Context Calibration (DC) [7], and Prototypical Calibration (PC) [10]. We reproduce all baselines strictly follows the authors’ instructions and recommendations to ensure a fair comparison.
Models and implementation details   We evaluate our method on Llama-2 7b and Llama-2 13b models [21]. For all experiments, unless stated otherwise, we use 1-shot ICL setting, i.e. one example per class, and repeat five times under different random seeds. We use $k=20$ sampes per class as the support set to obtain all threshold values by grid searching, as mentioned in the method section. The prompt template and more implementation details are specified in Appendix A.
@@
Figure 5: The performance comparison under different numbers of ICL shots.
@@ 4.2 Main Experiments
Table 1 presents the performance of various datasets and model sizes under the 1-shot setting. Our proposed UniBias method consistently achieves the highest accuracies in most cases. In terms of overall average accuracy, UniBias improves upon the standard ICL by a substantial margin of 3.39% and exceeds the state-of-the-art (SOTA) DC by 1.76% using Llama-2 7b. With Llama-2 13b, UniBias surpasses the standard ICL and the SOTA CC by 2.97% and 2.45%, respectively. Figure 5 further illustrates the results under zero-shot and various few-shot settings for COPA, SST2, and MMLU. Our proposed UniBias consistently surpasses other baselines in all scenarios, underscoring its effectiveness.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated impressive capabilities in various tasks using the in-context learning (ICL) paradigm. However, their effectiveness is often compromised by inherent bias, leading to prompt brittleness—sensitivity to design settings such as example selection, order, and prompt formatting. Previous studies have addressed LLM bias through external adjustment of model outputs, but the internal mechanisms that lead to such bias remain unexplored. Our work delves into these mechanisms, particularly investigating how feedforward neural networks (FFNs) and attention heads result in the bias of LLMs. By Interpreting the contribution of individual FFN vectors and attention heads, we identify the biased LLM components that skew LLMs’ prediction toward specific labels. To mitigate these biases, we introduce UniBias, an inference-only method that effectively identifies and eliminates biased FFN vectors and attention heads. Extensive experiments across 12 NLP datasets demonstrate that UniBias significantly enhances ICL performance and alleviates prompt brittleness of LLMs. The code is available at https://github.com/hzzhou01/UniBias.
@@ 1 Introduction
Large language models (LLMs) have shown exceptional capabilities in various natural language processing (NLP) tasks, employing the in-context learning (ICL) paradigm. This paradigm conditions LLMs on a context prompt comprising of a few example-label pairs (Brown et al., 2020; Wei et al., 2022; Dong et al., 2023; Zhou et al., 2024).
@@
Despite their impressive performance, LLMs are prone to prompt brittleness, characterized by high sensitivity to the choice (Zhao et al., 2021) and order (Lu et al., 2022) of examples, and prompt formatting (Min et al., 2022), as demonstrated in Figure 1. Such prompt brittleness is found to be arise from the bias in LLMs towards predicting certain answers (Zhao et al., 2021). The presence of the LLM bias undermines the robustness and adaptability of LLMs in diverse applications.
@@
Extensive research has focused on identifying factors that lead to LLM bias and strategies for mitigation. For instance, vanilla label bias (Fei et al., 2023) and recency bias (Zhao et al., 2021) demonstrate the LLM’s inherent non-contextual preference for certain labels and contextual preference for specific positions, respectively. Additionally, several calibration methods (Fei et al., 2023; Han et al., 2023; Zhao et al., 2021) are proposed to counteract the bias by adjusting decision boundaries of model output probabilities. However, these approaches are derived from external observations or adjustments of LLM outputs, leaving the internal mechanisms within LLMs that cause such bias poorly understood.
@@
In this work, we investigate the internal mechanism of LLM bias, specifically how feedforward neural networks (FFNs) and attention heads contribute to such bias. Building on findings in mechanistic interpretability (Elhage et al., 2021; Dar et al., 2023), we assess the contribution of individual attention heads and FFN vectors [^1] FFN vector refers to the value vector in the second weight matrix of the FFN layer. We elaborate on this in Section 2.1 to label predictions in LLMs. By identifying FFN vectors and attention heads that convey biased influences towards label prediction, we reveal the internal mechanisms behind several key bias factors, including vanilla label bias (Fei et al., 2023), recency bias (Zhao et al., 2021), and selection bias (Zheng et al., 2023). For instance, our analysis of FFN vectors without input context demonstrates that their cumulative impact biases the LLM towards specific labels, indicating a non-contextual preference for certain labels, i.e., vanilla label bias. We elaborate on the background of mechanistic interpretability in Section 2.1 and present our findings on the internal mechanisms of LLM biases in next section.
@@
Given our findings that various bias factors stem from the biased behaviors of attention heads and FFN vectors, we are prompted to ask: Can we identify the biased components of LLMs and mitigate their detrimental impact on label prediction? Motivated by this intuition, we propose UniBias, an inference-only method designed to identify and eliminate biased FFN vectors and attention heads in LLMs. Specifically, we begin by projecting each FFN vector and attention head into the vocabulary space to interpret the information conveyed by their outputs. We then detect biased components based on three criteria we defined: the relatedness criterion, the bias criterion, and the low variance criterion. After identification, we mitigate their impact by masking these biased components. Extensive experimental results demonstrate that LLMs, from which biased components have been removed, consistently outperform their original counterparts by a significant margin. Further, as illustrated in Figure 1, our method significantly improves both the performance and robustness of ICL with perturbations of various design settings.
@@
Figure 1: illustrates the prompt brittleness of ICL and the effectiveness of our method in mitigating this issue. Experiments are conducted in one-shot setting, using SST2 (Socher et al., 2013) dataset for experiments on example selection and prompt formatting and AGnews (Zhang et al., 2015) dataset for example order experiment due to more diverse combination of orders.
@@
  • In contrast to existing works based on external adjustments of LLM outputs, we mitigate LLM bias through manipulation of LLM internal structure. This novel perspective potentially offers a new direction for the field. Moreover, our method demonstrate an effective way to manipulate internal structures of LLMs.
  • We conduct a thorough investigation of the internal mechanisms underlying biases in LLMs, revealing the inner causes of these biases.
  • Extensive experiments across 12 NLP datasets demonstrate that, by removing the biased components, our UniBias method significantly enhances ICL performance and achieve state-of-the-art results. Additionally, it effectively addresses the issue of prompt brittleness.
@@ 2.1 Background
The theoretical background of this work is based on research on mechanistic interpretability (Elhage et al., 2021; Wang et al., 2022; Geva et al., 2021), which aims to explain the internal processes in language models (LMs), facilitating the interpretation of the contributions of individual model components to the final prediction.
@@
The Residual Stream We interpret Transformers following the view of residual stream (Elhage et al., 2021; Dar et al., 2023). Due to the residdual connection of Transformers, each layer takes a hidden state as input, and adds information obtained by its attention layer and FFN layer to the hidden state through residual connection. In this sence, the hidden state is a residual stream passed along layers, and each attention layer and FFN layer contribute to the final prediction by adding information to the residual stream.
Attention Heads   Following Elhage et al. (2021); Dar et al. (2023), the output of each attention layer of LM can be computed as the sum of all its attention heads. Specifically, for $l$-th layer, the input is ${X^{l}}\in\mathbb{R}^{N\times d}$, and the attention layer is parameterized by four matrices $W_{Q}^{l}$, $W_{K}^{l}$, $W_{V}^{l}$, $W_{O}^{l}\in\mathbb{R}^{d\times d}$. The columns of each projection matrix and the rows of the output matrix can be split into $H$ parts: $W_{Q}^{\ell,j},W_{K}^{\ell,j},W_{V}^{\ell,j}\in\mathbb{R}^{d\times\frac{d}{H}}$ and $W_{O}^{\ell,j}\in\mathbb{R}^{\frac{d}{H}\times d}$, where $H$ is the number of attention heads. We then find that:
@@
FFN   In line with Geva et al. (2021, 2022), transformer FFN layers can be cast as linear combination of vectors. Specifically, for an input vector $\mathbf{x}^{\ell}\in\mathbb{R}^{d}$, FFN parameter matrices $\mathbf{K}^{\ell},\mathbf{V}^{\ell}\in\mathbb{R}^{d_{m}\times d}$, the FFN output can be derived as:
@@
Logit Lens   The logit lens (Nostalgebraist, 2020) is a technique that directly decode hidden states into the vocabulary space using the unembedding matrix of the LLM for interpretation. This approach has been validated in various studies as an efficient method for interpreting the weight matrix or hidden states of LLMs (Dar et al., 2023; Hanna et al., 2023; Feng et al., 2024; Yu et al., 2023; Geva et al., 2021).
@@ 2.2 Internal Mechanisms of Bias Factors
We delve into the mechanisms behind several bias factors, analyzing the contributions of attention heads and FFN vectors to the biased predictions in LLMs. We explore vanilla label bias, position bias, and selection bias using the Llama-2 7B model (Touvron et al., 2023).
Vanilla Label Bias   The vanilla label bias (Fei et al., 2023), also known as common token bias (Zhao et al., 2021), is the inherent, uncontextual preference of the model towards predicting certain label names. Given the contextual nature of attention layers, our investigation focuses on the FFN layers, where we identified a corresponding uncontextual preference. Specifically, by projecting the FFN value vectors into the vocabulary space, we compute the logits for various label names for each FFN vector. Utilizing the residual stream insight, we then aggregate these logits for all FFN vectors whose label logits rank within the top $10$ over the vocabulary, reflecting uncontextual influences of FFN vectors that are effective in label prediction. This process yields what we term uncontextual accumulated FFN logits, revealing the intrinsic bias of the LLM towards predicting label names without the influence of input.
@@
Recency Bias   Recency bias refers to the tendency of LLMs to favor the label of the example at the end of the prompt (Zhao et al., 2021). By examining the behavior of attention heads within LLMs, we observe that specific heads consistently prioritize the example at the end of the prompt, providing an internal perspective on the origin of recency bias.
@@
Selection Bias   The selection bias refers that LLMs prefer to select specific option ID (like "Option A") as answers for multiple choice questions (Zheng et al., 2023). We have identified both FFN vectors and attention heads that consistently favor a specific option regardless of the ground truth label of the test sample, revealing the internal mechanism of selection bias.
We evaluate the Llama-2 7B model on the ARC dataset, which contains four options (A, B, C, D). We use a zero-shot setting to avoid the influence of position bias from multiple examples. More details are provided in Appendix A. Experimental results are illustrated in Figure 4. Firstly, we observe that the LLM exhibits a vanilla label bias favoring option "A", as shown in the first subfigure. Additionally, we identify a biased attention head that demonstrates a position bias consistently favoring the first option regardless of the ground truth labels of the test samples (second subfigure) or changes in the sequence of options (third subfigure). Since option A is usually the first option, these two biases both lead to the LLM’s preference for option A.
@@ 3.1 Biased FFN Vectors Identification
We use $p$ unlabeled samples from the task to assess the three criteria we defined. The coefficients and label logits of an FFN vector for these samples are denoted as $\mathbf{m}=[m_{0},m_{1},\ldots,\ m_{p-1}]$ and $\mathbf{G}=[\mathbf{g}^{(0)},\mathbf{g}^{(1)},\ldots,\mathbf{g}^{(p-1)}]^{\top}\in\mathbb{R}^{p\times c}$, respectively. An FFN vector is considered biased if it meets the following conditions, each corresponding to one of the three criteria we defined:
$\frac{1}{p}\sum_{k=0}^{p-1}\text{Sum}\left(\mathbf{G}_{k,:}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\text{Sum}\left(\mathbf{g^{(k)}}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\sum_{j=0}^{c-1}g_{j}^{(k)}>th_{FFN}^{1}$ (1)
$\frac{1}{p}\sum_{k=0}^{p-1}\text{Bias}\left(\mathbf{G}_{k,:}\right)=\frac{1}{p}\sum_{k=0}^{p-1}\text{Bias}\left(\mathbf{g^{(k)}}\right)=\frac{1}{p}\frac{1}{c}\sum_{k=0}^{p-1}\sum_{j=0}^{c-1}\left(g_{j}^{(k)}-\mu(\mathbf{g^{(k)}})\right)>th_{FFN}^{2}$ (2)
$CV\left(\mathbf{m}\right)=\frac{\sigma(\mathbf{m})}{\mu(\mathbf{m})}=\frac{\sqrt{\frac{1}{p}\sum_{j=0}^{p-1}\left(m_{k}-\mu(\mathbf{m})\right)^{2}}}{\frac{1}{p}\sum_{k=0}^{p-1}m_{k}}<th_{FFN}^{3}$ (3)
@@
The first equation corresponds to the relatedness criterion, measured by the sum of label logits. A higher sum indicates that the information introduced by the FFN vector is more relevant to label prediction. The second equation relates to the bias criterion, quantified by the deviation of the average logit for each label from the overall average logit across all labels. Ideally, for a set of test samples with different labels, the average logits for each label should be relatively balanced. A greater deviation from each label’s average compared to the overall average across all labels indicates a more biased distribution. The third equation addresses the low variance criterion, measured by the coefficient of variation (CV) of the FFN vector coefficients across different samples. The CV, calculated as the standard deviation normalized by the mean, indicates whether the label prediction information added by the FFN vector remains almost the same across different samples.
@@ 3.4 Grid Searching
Additionally, the support set can be replaced with unlabeled samples, using approximately twice the number of unlabeled samples compared to labeled ones. For further details, please see Appendix F.
@@ 4.1 Experimental Setup
Baselines   In addition to the standard ICL, we compare our proposed UniBias with state-of-the-art LLM debiasing and calibration baselines, including Contextual Calibration (CC) (Zhao et al., 2021), Domain-Context Calibration (DC) (Fei et al., 2023), and Prototypical Calibration (PC) (Han et al., 2023). We reproduce all baselines strictly follows the authors’ instructions and recommendations to ensure a fair comparison.
Models and implementation details   We evaluate our method using a range of LLMs, including Llama-2 7b, Llama-2 13b (Touvron et al., 2023), GPT-J (Wang and Komatsuzaki, 2021) and GPT2-XL (Radford et al., 2019). For all experiments, unless stated otherwise, we use 1-shot ICL setting, i.e. one example per class, and repeat five times under different random seeds. We use $k=20$ sampes per class as the support set to obtain all threshold values by grid searching, as mentioned in the method section. The prompt template and more implementation details are specified in Appendix A.
@@ 4.2 Main Experiments
Table 1 presents the performance of various datasets and model sizes under the 1-shot setting. Our proposed UniBias method consistently achieves the highest accuracies in most cases. In terms of overall average accuracy, UniBias improves upon the standard ICL by a substantial margin of 3.39% and exceeds the state-of-the-art (SOTA) DC by 1.76% using Llama-2 7b. With Llama-2 13b, UniBias surpasses the standard ICL and the SOTA CC by 2.97% and 2.45%, respectively. Figure 5 further illustrates the results under zero-shot and various few-shot settings for COPA, SST2, and MMLU. Additionally, to demonstrate the effectiveness of our method across different large language models, we present results for GPT-J and GPT2-XL in Figure 7 of Appendix C. Our proposed UniBias consistently surpasses other baselines in all scenarios, underscoring its effectiveness.
@@
Figure 5: The performance comparison under different numbers of ICL shots using Llama-2-7b.